OpenStudy (anonymous):

what is the slope of the line passing through the points (2,7) and (1-,3)?

OpenStudy (anonymous):

Can someone explain i keep getting 4 over 1

OpenStudy (yttrium):

m = (y2 - y1)/(x2-x1) m = (-3-7)/(1-2) m = -10/-1 m = 10

OpenStudy (anonymous):

2 over 7 3 over 4 4 over 3 1 over 3 are the possible answers how would i simplify 10?

OpenStudy (yttrium):

Where did you get that? No other answer rather than 10.

OpenStudy (jack1):

is that 2nd point supposed to be (-1, 3)?

OpenStudy (jack1):

@Yttrium u wanna redo?

OpenStudy (yttrium):

yea

OpenStudy (yttrium):

it should be m = (3-7)/(-1-2) m = -4/-3 m = 4/3

OpenStudy (anonymous):

what is the correct equation of the line shown below?

OpenStudy (yttrium):

m = (3-(-3))/(0-(-4)) Try to evaluate that.

OpenStudy (anonymous):

- 3 over 2 x - 3?

OpenStudy (anonymous):

I don't really know how to evaluate it. but out of the 4 possible answers that seems like the best answer.

OpenStudy (yttrium):

m = 6/4 or 3/2

OpenStudy (yttrium):

How did arrive that? Simple, just let P1 to be (-4,-3) and P2(0,3)\ then \[m = \frac{ 3-(-3) }{ 0-(-4) } = \frac{ 6 }{ 4 } = \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

these are the possible choices. So after i find out that it's 3 over 2 how do i find the rest out

OpenStudy (yttrium):

That's already the equation e. Okay. Using point (0,3), y - 3 = 3(x-0)/2 y = (3x/2)+3 --> A

OpenStudy (anonymous):

The x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line?

OpenStudy (anonymous):

I would think 5 over 2 but then i wouldn't know how to do the rest

OpenStudy (anonymous):

so how do i know if it is x - 5 or x - 2

OpenStudy (yttrium):

y-0 = 5(x+5)/2 y = (5x/2) + 5

OpenStudy (anonymous):

Last one What is the slope-intercept form of te equation y-7 = -5/\[y-7 =- \frac{ 5 }{ 2}(x+4)?\]

OpenStudy (anonymous):

just solve it , open the brackets , etc